OpenStudy (abbyrose0105):

Simplify: x^1/4 * x^5/4 * x^9/4 * 4√x

OpenStudy (imstuck):

Does it look like this:

OpenStudy (imstuck):

\[x ^{\frac{ 1 }{4 }}*x ^{\frac{ 5 }{ 4 }}*x ^{\frac{ 9 }{ 4 }}*4\sqrt{x}\]

OpenStudy (midhun.madhu1987):

 solve this